Mendel found that yellow pea pod color (G) was dominant to green pea pod color (g), and that round seeds (W) were dominant to wrinkled seeds (w). What percentage of each of the phenotypes below are expected from the following cross? GGww x GgWw

_____ Yellow and round

_____ Yellow and wrinkled

_____ Green and round

_____ Green and wrinkled

A. 0%

B. 25%

C. 33%

D. 50%

E. 66%

F. 75%

G. 100%
Answer #1

Yellow = GG, Gg

Green = gg

Round = WW, Ww

Wrinkled = ww

GGww × GgWw

Gametes Gw
GW GGWw yellow round
Gw GGww yellow wrinkled
gW GgWw yellow round
gw Ggww yellow wrinkled

Percentage of progenies -

Yellow round = 2/4 = 50%

Yellow wrinkled = 2/4 = 50%

Green round = 0%

Green wrinkled = 0%

Please rate high.